Paul Sabin organized Sabin Electronics 10 years ago to produce and sell several electronic devices on which he had secured paCost of goods sold Gross margin Selling and administrative expenses Net operating income Interest expense Net income before td. The return on equity. (Stockholders equity at the beginning of last year was $1,904,000.) e. Is the companys financial lYou decide next to assess the companys profitability. Compute the following for both this year and last year: a. The gross m
Answer #1

1.
a. Earning per share = Net Income / Common Stock Outstanding
This Year = 786800 / 44500 = $17.68 per share
Last Year = 422800 / 44500 = $9.50 per share
Common Stock outstanding = $890000 / 20 = 44500

b. Dividend Yield Ratio = Dividend per share / Market price x 100
This Year = $2.92 / 60 x 100 = 4.87%
Dividend per share = $130000/44500 = $2.92 per share
Last Year = $2.45 / 50 x 100 = 4.90%
Dividend per share = $109000/44500 = $2.45 per share

c. Dividend Payout ratio = Dividend / Net Income x 100
This Year = $130000 / 786800 = 16.52%
Last Year = $109000 / 422800 = 25.78%

d. Price Earning Ratio = Market price / EPS
This Year = $60 / 17.68 = 3.39 times
Last Year = $50 / 9.50 = 5.26 times

e. Book Value per share = Total Stockholder's Equity / Common Stock Outstanding
This Year = $2570800/44500 = $57.77 per share
Last Year = $1914000/44500 = $43.01 per share

As per HOMEWORKLIB RULES, we are supposed to answer 4 sub parts, i have done 5, so kindly post other part separately